# Tomur: Traffic-Aware Performance Prediction of On-NIC Network Functions with Multi-Resource Contention

### v3

@@ Abstract
Network function (NF) offloading on SmartNICs has been widely used in modern data centers, offering benefits in host resource saving and programmability. Co-running NFs on the same SmartNICs can cause performance interference due to onboard resource contention. Therefore, to meet performance SLAs while ensuring efficient resource management, operators need mechanisms to predict NF performance under such contention. However, existing solutions lack SmartNIC-specific knowledge and exhibit limited traffic awareness, leading to poor accuracy for on-NIC NFs. This paper proposes Tomur, a novel performance predictive system for on-NIC NFs. Tomur builds upon the key observation that co-located NFs contend for multiple resources, including onboard accelerators and the memory subsystem. It also facilitates traffic awareness according to the behaviors of individual resources to maintain accuracy as the external traffic attributes vary. Evaluation using BlueField-2 SmartNIC shows that Tomur improves the prediction accuracy by 78.8% and reduces SLA violations by 92.2% compared to state-of-the-art approaches, and enables new practical usecases.
Shaofeng Wu and Qiang Su contributed equally.
1 Introduction
Network Function Description Accelerator T Framework
FlowStats [42, 49] Flow statistics with aging. None Click
IPRouter [42, 49] L3 packet routing program. None Click
IPTunnel [29] L3 packet fragmentation tunnel. None Click
NAT [41, 34, 49] Network address translation for IPv4 based on MazuNAT. None Click
FlowMonitor [41, 47] Per-flow status and hardware payload scanning monitor. regex Click
NIDS [41, 42] Network intrusion prevention by hardware-based packet inspection. regex Click
IPComp Gateway [2, 54] Hardware-based payload scanning and compression. regex, compression Click
ACL [34, 17] Access control list based on DPDK ACL. None DPDK
FlowClassifier [17, 60] Flow tracking using hash table. None DPDK
FlowTracker [6] Packet processing pipeline with hardware-offloaded flow tracking. None DOCA
PacketFilter [6] A hardware-based pattern matching and PacketFilter. regex DOCA
Table 1: Typical NFs and accelerators they require from SmartNICs. Common resources (CPU, memory, and NIC subsystems) are not shown. T means that the NF performance heavily depends on the traffic attributes. The regex-based NFs use the same rule set from [8]. The last column indicates the programming framework we use to implement each NF.
@@ 1 Introduction
SmartNICs have been prevalent in modern data centers to deploy diverse network functions (NF) due to their benefits in programmability and host resource saving [49, 37, 57, 39, 50, 32, 41, 36]. They typically integrate heterogeneous onboard resources, such as SoC cores and domain-specific hardware accelerators, to cater to various NF demands. Moreover, vendors are developing increasingly resourceful NICs to meet evolving offloading needs [11, 55, 13]. To fully leverage these resources, it is common practice to co-locate multiple NFs on the same NIC [41, 53, 31, 36].
@@
Unfortunately, sharing SmartNIC resources among co-located NFs may lead to contention and performance degradation, posing challenges in maintaining the SLAs. One potential solution is to implement stringent resource isolation mechanisms on SmartNICs. Yet, previous work on this front requires specific NIC hardware architecture support and substantial rewriting of NF programs to accommodate new isolation abstractions [31], thus limiting their practical deployment. Consequently, developers still need extensive hand-tuning to ensure simultaneous SLA fulfillment for co-located NFs, which is time-consuming and error-prone [41, 31, 35].
@@
Ideally, if operators can predict the performance drop an NF will suffer before actually co-running it with other NFs on the same NIC, they can make better placement decisions on existing infrastructure and NF implementations. To achieve this, we need a systematic on-NIC NF performance prediction framework, which entails two new challenges.
@@
First, on-NIC NFs often utilize diverse onboard resources including memory and various hardware accelerators, making it common that contentions occur across heterogeneous resources. Prior work on NF performance prediction has primarily focused on memory subsystem contention as the sole source of performance interference for on-server NFs [42, 44, 28] only. We showcase that the state-of-the-art SLOMO [42] encounters high prediction errors when co-locating NFs contend for both memory and regex accelerator on a BlueField-2 SmartNIC, with $\sim$20in the median and $\sim$60in the worst case (2.2). The community lacks a clear understanding of (1) the impact of contention on individual domain-specific accelerators and (2) the overall effect of multi-resource contention taken together on performance.
@@
Second, NF performance is heavily influenced by traffic attributes such as flow counts and payload features, which changes dynamically for each NF. Current frameworks often either assume fixed traffic attributes [32, 49], or can only deal with a limited range of variations in these attributes (e.g., 20% in flow counts in [42]).
@@
In this paper, we propose Tomur, a new performance prediction framework that explicitly considers multi-resource contention and dynamic traffic attributes. Tomur conducts offline profiling of on-NIC NFs to collect their performance under diverse synthetic contention levels and traffic attributes. Leveraging these profiles, Tomur trains a contention-aware model for each NF, which is then used to predict the NF’s performance before its deployment, facilitating placement and other management decisions. We build Tomur for SoC SmartNICs due to their ease of programmability (e.g., DPDK and Click support), and tackle the above technical challenges by leveraging critical characteristics of on-NIC NFs.
@@
Multi-resource contention modeling. Tomur’s key idea here is to independently model individual resource contention and integrate these per-resource models together. We identify hardware accelerators and memory subsystems as primary sources of contention for on-NIC NFs. For accelerators, we find that it is a common design for NFs to interact with them through their own queues which are then coordinated by simple round-robin scheduling. This inspires us to take a white-box approach and propose a queueing-based contention model for accelerators. Memory subsystem contention can be modeled using a black-box ensemble-based ML model following existing work [42]. Then, to capture the end-to-end effect of each resource contention, we introduce execution-pattern-based composition. This makes intuitive sense because how each resource and its contention affects the overall performance critically depends on the NF’s execution pattern, e.g., pipeline or run-to-completion.
@@
Traffic-aware modeling. On top of multi-resource contention, Tomur employs traffic-aware augmentation to integrate knowledge of traffic attributes into per-resource models. Generally speaking, this can be done by feeding traffic attributes, e.g., flow count and packet size, as additional input features to per-resource models. Specifically, for accelerators, we can leverage the white-box nature of the queueing-based model and represent key model parameters as a function of traffic attributes; for memory subsystem which has a blackbox model, we simply fuse traffic attributes with performance counters as the input feature to extend the model. In addition, to curb the high profiling cost caused by the introduction of traffic attributes especially for black-box memory models, Tomur adopts adaptive profiling to prune traffic attribute dimensions and enforce targeted sampling at performance-critical ranges of the attributes.
@@
We implement Tomur in C and Python, leveraging typical offline profiling tools [19, 22, 18, 20] and sklearn [21], and evaluate it on 9 common NFs using the popular BlueField-2 SmartNIC. Our code is open source anonymously at [14]. Our testbed evaluation shows that Tomur achieves accurate NF throughput predictions under multi-resource contention and varying traffic attributes, with an average error of 3.7% across NFs which corresponds to 78.8% accuracy improvement compared to state-of-the-art SLOMO. As new usecases, we also illustrate that in NF placement, Tomur can reduce SLA violations by 88.5% and 92.2% compared to the classical greedy-based approaches [41, 53] and SLOMO, and in performance diagnosis it can deliver higher accuracy in identifying bottlenecks for on-NIC NFs.
2 Background and Motivation
@@
We start by presenting the brief background of network functions and their resource contention on SmartNICs, followed by the unique challenges of developing a contention-aware performance prediction framework which motivate our work.
2.1 Background
@@
SmartNICs have been widely used to offload various network functions (NFs) in modern data centers, mainly for their benefits in host resource saving and energy efficiency [49, 41, 37, 30, 53, 36]. The NFs leverage the onboard domain-specific hardware accelerators to achieve high throughput and low latency [41, 34, 42, 49]. We showcase some typical NFs seen across prior work [41, 34, 42, 49] and the types of resources they need in Table 1. Here Flow Monitor, NIDS, and PacketFilter require the regex accelerator for packet inspection and payload scanning, and IPComp Gateway requires both the regex and compression accelerators.
@@
Contention degrades performance Recently, co-running multiple NFs on the same SmartNIC has become more common to improve utilization [41, 53, 31, 36]. This can lead to performance degradation due to contention for shared resources. To demonstrate this effect, we profile the throughput drop of 9 typical NFs from Table 1 when they co-locate with other NFs. For each target NF, up to three other NFs are randomly selected (from Table 1). [^2] Some NFs require minimum two cores, while one BlueField-2 has eight cores in total. Each NF is given two dedicated cores while sharing the memory subsystem and hardware accelerators due to the lack of hardware- or system-level isolation support on current SmartNICs. Traffic profiles for all NFs consisted of 16K flows of 1500B packets, with flow sizes following the uniform distribution (further details in 7.1). For NFs processing payloads with regular expressions, we use exrex [15] to generate packet payloads. Note the packet arrival rates are set sufficiently high for all NFs to ensure it is not causing throughput drop. We measure the throughput drop ratio against the baseline when the target NF runs alone with two CPU cores, the entire memory and hardware accelerators. Figure 1 depicts the statistics of throughput drop ratios. We can see that when co-running with different (numbers and combinations of) NFs, resource contention can cause 4.2% to 62.2% throughput drop at the 95%ile, and 1.9% to 10.6% at the median.
@@
Figure 1: Throughput drop ratios of some NFs from Table 1 under resource contention when co-located with at most 3 other random NFs.
@@
New challenges. Modeling and predicting NF performance under resource contention is therefore of paramount importance for many management tasks [42, 32, 44, 28], and some prior work [42, 44, 28] has investigated this problem in network function virtualization where NFs run on commodity servers. An immediate question is, what makes contention-aware performance prediction different in the context of SmartNICs? We now highlight two unique challenges which are not well addressed in past efforts. Note all experiments in this section use the BlueField-2 (BF-2) SmartNIC.
2.2 Challenge 1: Multi-Resource Contention
Figure 2: Prediction errors (absolute percentage error) of FlowMonitor’s throughput using single-resource models. (a) Box and whisker plot of using only the memory-based SLOMO model or a regex-based model (4.1.1). We show the median error on top of each error box. (b) Mean average percentage error (MAPE) of sum and min composition of single-resource models. NF1 and NF2 adopt run-to-completion and pipeline resource usage pattern respectively.
@@ 2.2 Challenge 1: Multi-Resource Contention
We’ve seen that NFs on SmartNIC utilizes multiple heterogeneous onboard resources. Prior work, however, has only considered contention of the memory subsystems [42, 44, 28], missing the contention on other hardware accelerators. Their effectiveness as a result is tainted in the context of SmartNICs.
@@
To empirically substantiate our argument, we co-run FlowMonitor with up to three competing NFs chosen randomly from Table 1 on one BF-2. The traffic profiles are identical to the one in Figure 1. We use SLOMO [42] as the state-of-the-art memory-based prediction model and develop a new model for the regex accelerator due to lack of existing models (details in 4.1.1).
@@
We first train our single-resource models for FlowMonitor which uses regex accelerator in addition to CPU and memory, and validate their effectiveness under single-resource contention. We build two synthetic NFs, mem-bench and regex-bench, to assert controllable memory and regex contention, respectively, for generating training data (details in 6). Following SLOMO, we also collect data from our BF-2’s performance counters at runtime (e.g., memory read/write rates) as the model input. Absolute percentage error against FlowMonitor’s true throughput under single-resource contention is used as the loss function. Our models achieve the same $<$10% average prediction error for memory- and regex-only contention as reported in the SLOMO paper [42] (Tables 11 and 12 in Appendix A).
@@
In addition, NFs exhibit diverse execution patterns when utilizing these resources. For example, one NF may run in a pipeline manner for high throughput, while another may wait for the completion of dispatched requests to ensure low average latency (run-to-completion) [32, 20]. This makes composition of single-resource models, a strawman solution for multi-resource prediction, inaccurate.
To explore this, we analyze two simple composition approaches: (1) sum composition, which adds up the predicted throughput loss from each model [59, 32], and (2) min composition, which uses the minimal predicted throughput loss as the final output [41, 52]. Figure 2(b) presents the results of these two approaches in the same setting as Figure 2(a). We observe that while composition models reduce error, they do not guarantee optimal accuracy across all NFs. For NF1 with run-to-completion, sum composition works better, but its error is significant ($\sim$17%) for the pipeline NF2. The key reason is that the resource contention impact on end-to-end throughput varies by NF execution patterns. In pipeline-based NFs, throughput is constrained by the slowest stage on which resource contention causes the most significant performance interference compared. In contrast, for run-to-completion NFs, contention on different resources uniformly impacts the end-to-end throughput.
@@
2.3 Challenge 2: Traffic Attributes
An NF’s performance also depends on certain traffic attributes, such as number of flows, payload characteristics, etc., in many cases [32, 42, 49]. To see this, we measure FlowStat’s throughput when co-located with mem-bench, and vary mem-bench’s cache access rates (CAR). Figure 3(a) shows that FlowStat’s throughput drops differently in different traffic profiles as mem-bench’s CAR increases, implying that a traffic-agnostic model inevitably leads to high prediction errors when adapting to new traffic profiles.

### v5

# Performance Prediction of On-NIC Network Functions with Multi-Resource Contention and Traffic Awareness

@@ Abstract.
Network function (NF) offloading on SmartNICs has been widely used in modern data centers, offering benefits in host resource saving and programmability. Co-running NFs on the same SmartNICs can cause performance interference due to contention of onboard resources. To meet performance SLAs while ensuring efficient resource management, operators need mechanisms to predict NF performance under such contention. However, existing solutions lack SmartNIC-specific knowledge and exhibit limited traffic awareness, leading to poor accuracy for on-NIC NFs.
This paper proposes Yala, a novel performance predictive system for on-NIC NFs. Yala builds upon the key observation that co-located NFs contend for multiple resources, including onboard accelerators and the memory subsystem. It also facilitates traffic awareness according to the behaviors of individual resources to maintain accuracy as the external traffic attributes vary. Evaluation using BlueField-2 SmartNICs shows that Yala improves the prediction accuracy by 78.8% and reduces SLA violations by 92.2% compared to state-of-the-art approaches, and enables new practical usecases.
Network Function, SmartNIC, Resource Contention, Performance Prediction
ccs: Networks Network performance modelingccs: Hardware Networking hardware
1. Introduction
@@
SmartNICs have been prevalent in modern data centers to deploy diverse network functions (NF) due to their benefits in programmability and host resource saving (clara21, ; clicknp16, ; offpath22, ; bf2perf21, ; moore97, ; lognic23, ; e319, ; uno17, ; pipedevice22, ). They typically integrate heterogeneous onboard resources, such as SoC cores and domain-specific hardware accelerators, to cater to various NF demands. Moreover, vendors are developing increasingly resourceful NICs to meet evolving offloading needs (bf31, ; ipu, ; pensando, ). To fully leverage these resources, it is common practice to co-locate multiple NFs on the same NIC (e319, ; meili23, ; fairnic20, ; uno17, ).
@@
Unfortunately, sharing SmartNIC resources among co-located NFs may lead to contention and performance degradation, posing challenges in maintaining the SLAs. One potential solution is to implement stringent resource isolation mechanisms on SmartNICs. Yet, previous work on this front requires specific NIC hardware architecture support and substantial rewriting of NF programs to accommodate new isolation abstractions (fairnic20, ; osmosis24, ), thus limiting their practical deployment. Consequently, developers still need extensive hand-tuning to ensure simultaneous SLA fulfillment for co-located NFs, which is time-consuming and error-prone (e319, ; fairnic20, ; picnic20, ).
@@
Ideally, if operators can predict the performance drop an NF will suffer before actually co-running it with other NFs on the same NIC, they can make better resource management decisions on existing infrastructure and NF implementations. Concretely, SmartNIC platforms (e319, ; uno17, ; exoplane23, ; meili23, ) and SmartNIC-assisted clouds (awslam, ; db23, ; googlecloud, ) can maximize NF co-locations in SmartNIC offloading while minimizing SLA violations, which correspond to lower total cost of ownership (TCO) for providers and better experience for tenants, and enjoy faster diagnosis and reasoning of on-NIC NF performance compared to slow manual analysis (clara21, ; lognic23, ). To achieve these, we need a systematic on-NIC NF performance prediction framework, which entails two new challenges.
Network Function Accelerator T Framework
FlowStats (slomo20, ; clara21, ) None Click
IPRouter (slomo20, ; clara21, ) None Click
IPTunnel (click2000, ) None Click
NAT (e319, ; exoplane23, ; clara21, ) None Click
FlowMonitor (e319, ; metron18, ) Regex Click
NIDS (e319, ; slomo20, ) Regex Click
IPComp Gateway (comprbf, ; nfp17, ) Regex, compression Click
ACL (exoplane23, ; dpdk_pl, ) None DPDK
FlowClassifier (dpdk_pl, ; opennetvm16, ) None DPDK
FlowTracker (doca1, ) None DOCA
PacketFilter (doca1, ) Regex DOCA
Table 1. Typical NFs and accelerators they require from SmartNICs. Common resources (CPU, memory, and NIC subsystems) are not shown. T means that the NF performance heavily depends on the traffic attributes. The regex-based NFs use the same rule set from (l7filter, ). The last column indicates the programming framework we use to implement each NF.
@@
First, on-NIC NFs often utilize diverse onboard resources including memory and various hardware accelerators, making it common that contentions occur across heterogeneous resources. Prior work on NF performance prediction has primarily focused on memory subsystem contention as the sole source of performance interference for on-server NFs (slomo20, ; bubbleup11, ; dobre12, ). We showcase that the state-of-the-art SLOMO (slomo20, ) encounters high prediction errors when co-locating NFs contend for both memory and regex accelerator on a BlueField-2 SmartNIC, with $\sim$20in the median and $\sim$60in the worst case (2.2.1). The community lacks a clear understanding of (1) the impact of contention on individual domain-specific accelerators and (2) the overall effect of multi-resource contention on performance.
@@
Second, NF performance is heavily influenced by traffic attributes such as flow counts and payload features, which change dynamically for each NF. Current frameworks often either assume fixed traffic attributes (lognic23, ; clara21, ), or can only deal with a limited range of variations in these attributes (e.g., 20% in flow counts in (slomo20, )).
@@
In this paper, we propose Yala, a new performance prediction framework that explicitly considers multi-resource contention and dynamic traffic attributes. Yala conducts offline profiling of on-NIC NFs to collect their performance under diverse synthetic contention levels and traffic attributes. Leveraging these profiles, Yala trains a contention- and traffic-aware model for each NF, which is then used to predict the NF’s performance before its deployment, facilitating placement and other management decisions. We build Yala for SoC SmartNICs due to their ease of programmability (e.g., DPDK and Click support), and tackle the above technical challenges by leveraging critical characteristics of on-NIC NFs.
@@
Multi-resource contention modeling. Yala’s key idea here is to independently model individual resource contention and integrate these per-resource models together. We identify hardware accelerators and memory subsystems as primary sources of contention for on-NIC NFs. For accelerators, we find that it is a common design for NFs to interact with them through their own queues which are coordinated by round-robin scheduling. This inspires us to take a white-box approach and propose a queueing-based contention model. Memory subsystem contention can be modeled using a black-box ensemble-based ML model following existing work (slomo20, ). Then, to capture the end-to-end effect of each resource, we introduce execution-pattern-based composition. This makes intuitive sense because how each resource and its contention affects the overall performance critically depends on whether NF runs as a pipeline or in a run-to-completion fashion.
@@
Traffic-aware modeling. On top of multi-resource contention, Yala employs traffic-aware augmentation to integrate knowledge of traffic attributes into per-resource models. Generally speaking, this can be done by feeding traffic attributes, e.g., flow count and packet size, as additional features to per-resource models. Specifically, for accelerators, we can leverage the white-box nature of the queueing-based model and represent key model parameters as a function of traffic attributes; for memory subsystem which has a blackbox model, we simply fuse traffic attributes with performance counters as features to extend the model. In addition, to curb the high profiling cost caused by the introduction of traffic attributes especially for black-box memory models, Yala adopts adaptive profiling to prune attribute dimensions and enforce targeted sampling at performance-critical ranges of the attributes.
@@
We implement Yala in C and Python, leveraging typical offline profiling tools (perf, ; stressng, ; mbw, ; RXPbench, ) and sklearn (sklearn, ), and evaluate it on 9 common NFs using BlueField-2 SmartNIC. Our code is open source anonymously at (yalagithub, ). Our testbed evaluation shows that Yala achieves accurate NF throughput predictions under multi-resource contention and varying traffic attributes, with an average error of 3.7% across NFs which corresponds to 78.8% improvements compared to state-of-the-art SLOMO. As new usecases, we also illustrate that in NF placement, Yala can reduce SLA violations by 88.5% and 92.2% compared to greedy approaches (e319, ; meili23, ) and SLOMO, and in performance diagnosis it can deliver higher accuracy in identifying bottlenecks for on-NIC NFs.
2. Background and Motivation
@@
We start by presenting the brief background of network function resource contention on SmartNICs, followed by the unique challenges of developing a contention-aware performance prediction framework.
2.1. Background
@@
SmartNICs have been widely used to offload various network functions (NFs) in modern data centers, mainly for their benefits in host resource saving and energy efficiency (clara21, ; e319, ; clicknp16, ; azure18, ; meili23, ; uno17, ). The NFs leverage the onboard domain-specific hardware accelerators to achieve high throughput and low latency (e319, ; exoplane23, ; slomo20, ; clara21, ). We showcase some typical NFs seen across prior work (e319, ; exoplane23, ; slomo20, ; clara21, ) and the types of resources they need in Table 1. Here Flow Monitor, NIDS, and PacketFilter require the regex accelerator for packet inspection and payload scanning, and IPComp Gateway requires both the regex and compression accelerators.
@@
Contention degrades performance Recently, co-running multiple NFs on the same SmartNIC has become more common to improve utilization (e319, ; meili23, ; fairnic20, ; uno17, ). This can lead to performance degradation due to contention for shared resources. To demonstrate this effect, we profile the throughput drop of 9 typical NFs from Table 1 when they co-locate with other NFs. For each target NF, up to three other NFs are randomly selected (from Table 1) [^1] Some NFs require minimum two cores, while one BlueField-2 has eight cores in total.. Each NF is given two dedicated cores while sharing the memory subsystem and hardware accelerators due to the lack of hardware- or system-level isolation support on current SmartNICs. Traffic profiles for all NFs consisted of 16K flows of 1500B packets, with flow sizes following the uniform distribution (further details in 7.1). For NFs processing payloads with regular expressions, we use exrex (exrex, ) to generate packet payloads. Note the packet arrival rates are set sufficiently high for all NFs to ensure it is not causing throughput drop. We measure the throughput drop ratio against the baseline when the target NF runs alone with two CPU cores, the entire memory and hardware accelerators. Figure 1 depicts the statistics of throughput drop ratios. We can see that when co-running with different (numbers and combinations of) NFs, resource contention can cause 4.2% to 62.2% throughput drop at the 95%ile, and 1.9% to 10.6% at the median.
@@
Figure 1. Throughput drop ratios of some NFs from Table 1 under resource contention when co-located with at most 3 other random NFs.
[…]
2.2. Challenges
@@
Modeling and predicting NF performance under resource contention is therefore of paramount importance for many management tasks (slomo20, ; lognic23, ; bubbleup11, ; dobre12, ), and some prior work (slomo20, ; bubbleup11, ; dobre12, ) has investigated this problem in network function virtualization where NFs run on commodity servers. An immediate question is, what makes contention-aware performance prediction different in the context of SmartNICs? We now highlight two unique challenges which are not well addressed in past efforts. Note all experiments in this section use the BlueField-2 (BF-2) SmartNIC.
2.2.1. Multi-Resource Contention
@@
We’ve seen that NFs on SmartNIC utilizes multiple heterogeneous onboard resources. Prior work, however, has only considered contention of the memory subsystems (slomo20, ; bubbleup11, ; dobre12, ), missing the contention on other hardware accelerators. Their effectiveness as a result is tainted in the context of SmartNICs.
@@
To empirically substantiate our argument, we co-run FlowMonitor with up to three competing NFs chosen randomly from Table 1 on one BF-2. The traffic profiles are identical to the one in Figure 1. We use SLOMO (slomo20, ) as the state-of-the-art memory-based prediction model and develop a new model for the regex accelerator due to lack of existing models (details in 4.1.1).
@@
We first train our single-resource models for FlowMonitor which uses regex accelerator in addition to CPU and memory, and validate their effectiveness under single-resource contention. We build two synthetic NFs, mem-bench and regex-bench  [^2] We build synthetic NFs for three main purposes: 1) collecting training data, 2) exploring insights that support our design choices, and 3) microbenchmarks. For example, regex-bench is purpose-built to have negligible memory subsystem usage but extensive regex accelerator usage, and we rely on it to investigate the contention behavior in regex accelerator (4.1.1). For evaluations on end-to-end accuracy (7.2, 7.3, 7.4) and use cases (7.5), we employ real NFs from Table 1 instead., to assert controllable memory and regex contention, respectively, for generating training data (details in 6). Following SLOMO, we also collect data from our BF-2’s performance counters at runtime (e.g., memory read/write rates) as the model input. Absolute percentage error against FlowMonitor’s true throughput under single-resource contention is used as the comparing metric. Our models achieve the same $<$10% average prediction error for memory- and regex-only contention as reported in the SLOMO paper (slomo20, ).
@@
In addition, NFs exhibit diverse execution patterns when utilizing these resources. For example, one NF may run in a pipeline manner for high throughput, while another may wait for the completion of dispatched requests to ensure low average latency (run-to-completion(lognic23, ; RXPbench, ). This makes composition of single-resource models, a strawman solution for multi-resource prediction, inaccurate.
To explore this, we analyze two simple composition approaches: (1) sum composition, which adds up the predicted throughput loss from each model (nnmeter, ; lognic23, ), and (2) min composition, which uses the maximum predicted throughput loss as the final output (e319, ; flextoe22, ). Figure 2(b) presents the results of these two approaches in the same setting as Figure 2(a). We observe that while composition models reduce error, they do not guarantee optimal accuracy across all NFs. For NF1 with run-to-completion, sum composition works better, but its error is significant ($\sim$17%) for the pipeline NF2. The key reason is that the resource contention impact on end-to-end throughput varies by NF execution patterns. In pipeline-based NFs, throughput is constrained by the slowest stage on which resource contention causes the most significant performance interference compared. In contrast, for run-to-completion NFs, contention on different resources uniformly impacts the end-to-end throughput.
@@
Figure 2. Prediction errors (absolute percentage error) of FlowMonitor’s throughput using single-resource models. (a) Box and whisker plot of using only the memory-based SLOMO model or a regex-based model (4.1.1). We show the median error on top of each error box. (b) Mean average percentage error (MAPE) of sum and min composition of single-resource models. NF1 and NF2 adopt run-to-completion and pipeline resource usage pattern respectively.
[…]
2.2.2. Traffic Attributes
An NF’s performance also depends on certain traffic attributes, such as number of flows, payload characteristics, etc., in many cases (lognic23, ; slomo20, ; clara21, ). To see this, we measure FlowStat’s throughput when co-located with mem-bench, and vary mem-bench’s cache access rates (CAR). Figure 3(a) shows that FlowStat’s throughput drops differently in different traffic profiles as mem-bench’s CAR increases, implying that a traffic-agnostic model inevitably leads to high prediction errors when adapting to new traffic profiles.